Cedric Company recently traded in an older model of equipment for a new model. The old model’s book value was $360,000 (original cost of $780,000 less $420,000 in accumulated depreciation) and its fair value was $400,000. Cedric paid $80,000 to complete the exchange which has commercial substance.
 
Required:
Prepare the journal entry to record the exchange. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1

dr. equipment- new                       480,000

dr. accumulated depreciation        420,000

  cr. gain on exchange of assets                   40,000

  cr. cash                                                        80,000

  cr. equipment-old                                        780,000


equipment 400,000+80,000= 480,000

gain= 400,000-360,000= 40,000

answered by: anonymous